DIRECT-ABSORPTION OBSERVATION OF A DOUBLY-FORBIDDEN ELECTRONIC TRANSITION IN DIACETYLENE USING CAVITY RING-DOWN SPECTROSCOPY

\underline{B. GILES}, C. A. ARRINGTON, F. C. HAGEMEISTER, AND T. S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907-1393.

The nominally ^{3}Δ_{u} ← X^{1}σ_{g}^{+} transition in room-temperature diacetylene has been observed by direct absorption for the first time using the cavity ring-down technique pioneered by O’Keefe and Deacon ^{a}. Using mirrors with high reflectivity (99.85%) the wavelength range 396 nm - 342 nm was investigated. Two bands at 27090 cm^{−1} and 27430 cm^{−1} are observed which show a red shading consistent with a π^{∗} ← π transition. Band contours also suggest a linear to bent electronic excitation. The largest spectral feature has a measured molar absorption coefficient of 1.5x10^{−4}Lmol^{−1}cm^{−1}. The ^{3}Δ_{u} state is one of the states thought to be responsible for photochemical reactivity.

aA. O’Keefe and D. A. G. Deacon, Rev. Sci. Instrum. 59, 2544 (1988).